OpenStudy (anonymous):

Solve: sqrt 4x+1 - sqrt x-2 = 3

OpenStudy (campbell_st):

square every term in the equation then collect like terms and then solve for x

OpenStudy (campbell_st):

use the rule \[(\sqrt{x})^2 = x\]

OpenStudy (anonymous):

is the answer supposed to be x=10/3?

jimthompson5910 (jim_thompson5910):

if the original problem is \[\Large \sqrt{4x+1}-\sqrt{x-2} = 3\] then there are 2 solutions and they are both whole numbers